# Accelerating Ill-conditioned Hankel Matrix RecoveryHanQin Cai, Longxiu Huang, Xiliang Lu, and Juntao You

### v1

@@ 5 Proofs
In this section, we provide the analysis of the claimed theoretical results. All the proofs are under Sections 1.1, 1.1 and 1.1. We start with introducing some additional notations used in the analysis.
For ease of presentation, in addition to Section 1.2, we denote the following slightly abused notation for the proofs. We let $\Omega$ be any subset of $[n]$ that satisfies Section 1.1 with $m:=\lvert\Omega\rvert$ and $p:=m/n$. We denote the tangent space of rank-$r$ matrix manifold at $\mathcal{G}\bm{z}_{\star}$ by
@@
(10) $\Omega_{a}:=\{(i,j):i+j=a+1,1\leq i\leq n_{1},1\leq j\leq n_{2}\},~{}a=1,2\cdots,n.$
We also introduce the following notations for the ease of presentation: Let ${\bm{Q}}_{k}$ be the best rotation matrix that aligns $({\bm{L}}_{k},{\bm{R}}_{k})$ and $(\bm{L}_{\star},\bm{R}_{\star})$, and according to [33, Lemma 22] we know ${\bm{Q}}_{k}$ always exist if $d_{k}<\sigma^{\star}_{r}$. Denote
$\bm{L}_{\natural}:={\bm{L}}_{k}{\bm{Q}}_{k},~{}\bm{R}_{\natural}:={\bm{R}}_{k}{\bm{Q}}_{k}^{-*},~{}\bm{\Delta}_{{\bm{L}}_{k}}:=\bm{L}_{\natural}-\bm{L}_{\star},~{}\bm{\Delta}_{{\bm{R}}_{k}}:=\bm{R}_{\natural}-\bm{R}_{\star}.$
@@ 5.1 Technical lemmas
Lemma 5.1 ([12, Lemma 2][4, Lemma 9]).
@@
(11) $\left\lVert\left(p^{-1}\mathcal{G}\Pi_{\Omega}-\mathcal{G}\right)\bm{z}_{\star}\right\lVert_{2}\leq\tilde{c}_{0}\sqrt{(\mu c_{s}r\log n)/m}\left\lVert\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$
@@
Proof 5.2.
The bound of $\left\lVert\left(p^{-1}\mathcal{G}\Pi_{\Omega}-\mathcal{G}\right)\bm{z}_{\star}\right\lVert_{2}$ is provided in [12, Lemma 2] (see [12, Eq. (25)]), and also proved with a slightly different sampling model in [4, Lemma 9].
Lemma 5.3 ([4, Lemma 10]).
There exists a universal constant $\tilde{c}_{1}$ such that if provided $m\geq\varepsilon_{0}^{-2}\tilde{c}_{1}\mu r\log n$, it holds
(12) $\left\lVert p^{-1}\mathcal{P}_{T}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}\mathcal{P}_{T}-\mathcal{P}_{T}\mathcal{G}\mathcal{G}^{*}\mathcal{P}_{T}\right\lVert\leq\varepsilon_{0}$
with probability at least $1-2n^{-2}$.
Proof 5.4.
The result is proved in [4, Lemma 9] under a slightly different sampling model and can be extended to our sampling model, which is omitted here for simplicity.
Lemma 5.5 ([11, Lemma 5]).
For all $\bm{u}\in\mathbb{R}^{n_{1}}$, $\bm{v}\in\mathbb{R}^{n_{2}}$, it holds
(13) $p^{-1}\sum_{a_{k}\in\Omega}\sum_{(i,j)\in\Omega_{a_{k}}}u_{i}v_{j}\leq\left\lVert\bm{u}\right\lVert_{1}\left\lVert\bm{v}\right\lVert_{1}+\sqrt{24np^{-1}\log n}\left\lVert\bm{u}\right\lVert_{2}\left\lVert\bm{v}\right\lVert_{2}$
@@
Lemma 5.6.
If $m\geq\log n$, then under event (13) it holds
$~{}p^{-1}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right),\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right\rangle\right\rvert$
$\leq$ $~{}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}+\sqrt{24p^{-1}n\log n}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}},$
@@
$~{}p^{-1}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right),\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right\rangle\right\rvert$
$=$ $~{}p^{-1}\sum_{k=1}^{m}\bigg{|}\sum_{(i,j)\in\Omega_{a_{k}}}\frac{1}{\sqrt{\varsigma_{a_{k}}}}\left\langle{\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*},\bm{e}_{i}\bm{e}_{j}^{\top}\right\rangle\bigg{|}^{2}$
$\leq$ $~{}p^{-1}\sum_{k=1}^{m}\sum_{(i,j)\in\Omega_{a_{k}}}\frac{1}{\sqrt{\varsigma_{a_{k}}}}\left\lVert{\bm{L}}_{{\bm{A}}}(i,:)\right\lVert_{2}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}(j,:)\right\lVert_{2}^{2}$
$\leq$ $~{}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}+\sqrt{24p^{-1}n\log n}\sqrt{\sum_{i}\left\lVert{\bm{L}}_{{\bm{A}}}(i,:)\right\lVert_{2}^{4}}\sqrt{\sum_{j}\left\lVert{\bm{R}}_{{\bm{A}}}(j,:)\right\lVert_{2}^{4}}$
$\leq$ $~{}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}+\sqrt{24p^{-1}n\log n}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}.$
where the second line follows from Cauchy-Schwarz inequality and the third line follows from Lemma 5.5.
Lemma 5.8.
For all matrices ${\bm{A}},{\bm{B}}\in T$, under event (12), it holds
$\left\lvert\left\langle\mathcal{G}(p^{-1}\Pi_{\Omega}-\mathcal{I})\mathcal{G}^{*}\bm{A},\bm{B}\right\rangle\right\rvert\leq\varepsilon_{0}\left\lVert{\bm{A}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{B}}\right\lVert_{\mathrm{F}},$
@@
Proof 5.9.
For any matrices ${\bm{A}},{\bm{B}}\in T$, we have $\mathcal{P}_{T}{\bm{A}}={\bm{A}}$ and $\mathcal{P}_{T}{\bm{B}}={\bm{B}}$. Then, under event (12), it holds
$\left\lvert\left\langle\mathcal{G}(p^{-1}\Pi_{\Omega}-\mathcal{I})\mathcal{G}^{*}{\bm{A}},{\bm{B}}\right\rangle\right\rvert=\left\lvert\left\langle\mathcal{P}_{T}\mathcal{G}(p^{-1}\Pi_{\Omega}-\mathcal{I})\mathcal{G}^{*}\mathcal{P}_{T}{\bm{A}},{\bm{B}}\right\rangle\right\rvert\leq\varepsilon_{0}\left\lVert{\bm{A}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{B}}\right\lVert_{\mathrm{F}},$
@@
(14) $\big{\|}\mathcal{W}\Pi_{\Omega}\bm{s}_{\star}-\mathcal{W}\bm{s}\big{\|}_{\infty}\leq 2\left\lVert\mathcal{W}\Pi_{\Omega}(\bm{z}_{\star}-\bm{z})\right\lVert_{\infty}.$
Proof 5.11.
Recall $\bm{f}=\bm{z}_{\star}+\bm{s}_{\star}$. By the definition of $\bm{s}$, we have
@@
Combing all the pieces we have $\big{\|}\mathcal{W}\Pi_{\Omega}\bm{s}_{\star}-\mathcal{W}\bm{s}\big{\|}_{\infty}\leq 2\left\lVert\mathcal{W}\Pi_{\Omega}(\bm{z}_{\star}-\bm{z})\right\lVert_{\infty}$.
Lemma 5.12.
@@
(15) $\left\lVert(\widetilde{{\bm{R}}}\widetilde{{\bm{Q}}}^{-*}-\bm{R}_{\star})\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}$ $\leq\left\lVert(\widetilde{{\bm{R}}}\widetilde{{\bm{Q}}}^{-*}-\bm{R}_{\star})\bm{\Sigma}_{\star}^{1/2}\bm{\Sigma}_{\star}^{-1}\right\lVert_{\mathrm{F}}$
(16) $\leq\left\lVert\bm{\Sigma}_{\star}^{-1}\right\lVert_{2}\mathrm{dist}\left(\widetilde{{\bm{L}}},\widetilde{{\bm{R}}};\bm{L}_{\star},\bm{R}_{\star}\right)\leq\varepsilon_{0}.$
@@
(17) $\frac{C}{\big{\|}\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\big{\|}_{2}}\geq\frac{C}{(1+\varepsilon_{0})\big{\|}\widetilde{{\bm{L}}}_{i,:}\widetilde{{\bm{Q}}}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}}\geq\frac{\big{\|}{\bm{L}}_{i,:}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}}{\big{\|}\widetilde{{\bm{L}}}_{i,:}\widetilde{{\bm{Q}}}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}},$
@@
For any vectors $\bm{u},\bm{v}\in\mathbb{R}^{n}$, it holds $\left\lVert\min(1,\lambda)\bm{u}-\bm{v}\right\lVert_{2}\leq\left\lVert\bm{u}-\bm{v}\right\lVert_{2}$ provided $\lambda\geq\frac{\left\lVert\bm{v}\right\lVert_{2}}{\left\lVert\bm{u}\right\lVert_{2}}$ (see [33, Claim 5]). Recalling that ${\bm{L}}_{i,:}=\min\left(1,\frac{C}{\big{\|}\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\big{\|}_{2}}\right)\widetilde{{\bm{L}}}_{i,:}$, thus by Eq. 17 we have
@@
and similarly $\left\lVert\left({\bm{R}}_{i,:}\widetilde{{\bm{Q}}}^{-*}-[\bm{R}_{\star}]_{i,:}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2}\leq\left\lVert\left(\widetilde{{\bm{R}}}_{i,:}\widetilde{{\bm{Q}}}^{-*}-[\bm{R}_{\star}]_{i,:}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2}$. Then, by definition of the metric we have the locally non-expensiveness of $\mathcal{P}_{\mathcal{C}}$, i.e.,
@@
$\leq\mathrm{dist}\left(\widetilde{{\bm{L}}},\widetilde{{\bm{R}}};\bm{L}_{\star},\bm{R}_{\star}\right).$
Noticing that $\|{\bm{L}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{1/2}\|_{2}=\|\min(1,\frac{C}{\|\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\|_{2}})\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{1/2}\|_{2}\leq C$, $\forall i\in[n_{1}]$, thus $\|{\bm{L}}{\bm{R}}^{*}\|_{2,\infty}\leq\|{\bm{L}}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{1/2}\|_{2,\infty}\|{\bm{R}}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{-1/2}\|_{2}\leq C$. Similarly, we have $\|{\bm{R}}{\bm{L}}^{*}\|_{2,\infty}\leq C$.
5.2 Proof of Proposition 3.1
Proof 5.14.
For $d_{k}\leq\varepsilon_{0}\sigma^{\star}_{r}$, following Eq. 15, we have $\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}\leq\varepsilon_{0},$ thus

### v2

@@ 5 Proofs
In this section, we provide the analysis of the claimed theoretical results. All the proofs are under Assumptions 1, 2 and 3. We start with introducing some additional notations used in the analysis.
For ease of presentation, in addition to Section 1.2, we denote the following slightly abused notation for the proofs. We let $\Omega$ be any subset of $[n]$ that satisfies Assumption 1 with $m:=\lvert\Omega\rvert$ and $p:=m/n$. We denote the tangent space of rank-$r$ matrix manifold at $\mathcal{G}\bm{z}_{\star}$ by
@@
$\Omega_{a}:=\{(i,j):i+j=a+1,1\leq i\leq n_{1},1\leq j\leq n_{2}\},~{}a=1,2\cdots,n.$ (8)
We also introduce the following notations for the ease of presentation: Let ${\bm{Q}}_{k}$ be the best rotation matrix that aligns $({\bm{L}}_{k},{\bm{R}}_{k})$ and $(\bm{L}_{\star},\bm{R}_{\star})$, and according to [36, Lemma 22] we know ${\bm{Q}}_{k}$ always exist if $d_{k}<\sigma^{\star}_{r}$. Denote
$\bm{L}_{\natural}:={\bm{L}}_{k}{\bm{Q}}_{k},~{}\bm{R}_{\natural}:={\bm{R}}_{k}{\bm{Q}}_{k}^{-*},~{}\bm{\Delta}_{{\bm{L}}_{k}}:=\bm{L}_{\natural}-\bm{L}_{\star},~{}\bm{\Delta}_{{\bm{R}}_{k}}:=\bm{R}_{\natural}-\bm{R}_{\star}.$ (9)
@@ 5.1 Technical lemmas
In this subsection, we mainly introduce some supporting lemmas, including bounds estimated from probability inequalities.
Lemma 1 ([12, Lemma 2][4, Lemma 9]).
@@
$\left\lVert\left(\frac{1}{p}\mathcal{G}\Pi_{\Omega}-\mathcal{G}\right)\bm{z}_{\star}\right\lVert_{2}\leq\tilde{c}_{0}\sqrt{(\mu c_{s}r\log n)/m}\left\lVert\mathcal{G}\bm{z}_{\star}\right\lVert_{2}$ (10)
@@
Proof.
The bound of $\left\lVert\left(\frac{1}{p}\mathcal{G}\Pi_{\Omega}-\mathcal{G}\right)\bm{z}_{\star}\right\lVert_{2}$ is provided in [12, Lemma 2] (see [12, Eq. (25)]), and also proved with a slightly different sampling model in [4, Lemma 9]. ∎
Lemma 2 ([4, Lemma 10]).
There exists a universal constant $\tilde{c}_{1}$ such that
$\left\lVert\frac{1}{p}\mathcal{P}_{T}\mathcal{G}\Pi_{\Omega}\mathcal{G}^{*}\mathcal{P}_{T}-\mathcal{P}_{T}\mathcal{G}\mathcal{G}^{*}\mathcal{P}_{T}\right\lVert\leq\varepsilon_{0}$ (11)
holds with probability at least $1-2n^{-2}$, provided $m\geq\varepsilon_{0}^{-2}\tilde{c}_{1}\mu r\log n$.
Proof.
The result is proved in [4, Lemma 9] under a slightly different sampling model and can be extended to our sampling model, which is omitted here for simplicity. ∎
Lemma 3 ([11, Lemma 5]).
For all $\bm{u}\in\mathbb{R}^{n_{1}}$, $\bm{v}\in\mathbb{R}^{n_{2}}$, it holds that
$\frac{1}{p}\sum_{a_{k}\in\Omega}\sum_{(i,j)\in\Omega_{a_{k}}}u_{i}v_{j}\leq\left\lVert\bm{u}\right\lVert_{1}\left\lVert\bm{v}\right\lVert_{1}+\sqrt{\frac{24n\log n}{p}}\left\lVert\bm{u}\right\lVert_{2}\left\lVert\bm{v}\right\lVert_{2}$ (12)
@@
Lemma 4.
If $m\geq\log n$, then under event (12) it holds that
$~{}\frac{1}{p}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right),\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right\rangle\right\rvert$
$\leq$ $~{}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}+\sqrt{\frac{24n\log n}{p}}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}},$
@@
$~{}\frac{1}{p}\left\lvert\left\langle\Pi_{\Omega}\mathcal{G}^{*}\left({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*}\right),\mathcal{G}^{*}({\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*})\right\rangle\right\rvert$
$=$ $~{}\frac{1}{p}\sum_{k=1}^{m}\bigg{|}\sum_{(i,j)\in\Omega_{a_{k}}}\frac{1}{\sqrt{\varsigma_{a_{k}}}}\left\langle{\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*},\bm{e}_{i}\bm{e}_{j}^{\top}\right\rangle\bigg{|}^{2}$
$\leq$ $~{}\frac{1}{p}\sum_{k=1}^{m}\sum_{(i,j)\in\Omega_{a_{k}}}\bigg{|}\left\langle{\bm{L}}_{{\bm{A}}}{\bm{R}}_{{\bm{A}}}^{*},\bm{e}_{i}\bm{e}_{j}^{\top}\right\rangle\bigg{|}^{2}$
$\leq$ $~{}\frac{1}{p}\sum_{k=1}^{m}\sum_{(i,j)\in\Omega_{a_{k}}}\left\lVert{\bm{L}}_{{\bm{A}}}(i,:)\right\lVert_{2}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}(j,:)\right\lVert_{2}^{2}$
$\leq$ $~{}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}+\sqrt{\frac{24n\log n}{p}}\sqrt{\sum_{i}\left\lVert{\bm{L}}_{{\bm{A}}}(i,:)\right\lVert_{2}^{4}}\sqrt{\sum_{j}\left\lVert{\bm{R}}_{{\bm{A}}}(j,:)\right\lVert_{2}^{4}}$
$\leq$ $~{}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}^{2}+\sqrt{\frac{24n\log n}{p}}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{2,\infty}\left\lVert{\bm{L}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{R}}_{{\bm{A}}}\right\lVert_{\mathrm{F}}.$
where the first and second inequalities follow from the Cauchy-Schwarz inequality, and the third inequality follows from Lemma 3. ∎
Lemma 5.
For all matrices ${\bm{A}},{\bm{B}}\in T$, under event (11), it holds that
$\left\lvert\left\langle\mathcal{G}(\frac{1}{p}\Pi_{\Omega}-\mathcal{I})\mathcal{G}^{*}\bm{A},\bm{B}\right\rangle\right\rvert\leq\varepsilon_{0}\left\lVert{\bm{A}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{B}}\right\lVert_{\mathrm{F}},$
@@
Proof.
For any matrices ${\bm{A}},{\bm{B}}\in T$, we have $\mathcal{P}_{T}{\bm{A}}={\bm{A}}$ and $\mathcal{P}_{T}{\bm{B}}={\bm{B}}$. Then, under event (11), it holds that
$\left\lvert\left\langle\mathcal{G}(\frac{1}{p}\Pi_{\Omega}-\mathcal{I})\mathcal{G}^{*}{\bm{A}},{\bm{B}}\right\rangle\right\rvert=\left\lvert\left\langle\mathcal{P}_{T}\mathcal{G}(\frac{1}{p}\Pi_{\Omega}-\mathcal{I})\mathcal{G}^{*}\mathcal{P}_{T}{\bm{A}},{\bm{B}}\right\rangle\right\rvert\leq\varepsilon_{0}\left\lVert{\bm{A}}\right\lVert_{\mathrm{F}}\left\lVert{\bm{B}}\right\lVert_{\mathrm{F}},$
@@
$\big{\|}\mathcal{W}\Pi_{\Omega}\bm{s}_{\star}-\mathcal{W}\bm{s}\big{\|}_{\infty}\leq 2\left\lVert\mathcal{W}\Pi_{\Omega}(\bm{z}_{\star}-\bm{z})\right\lVert_{\infty}.$ (13)
Proof.
Recall $\bm{f}=\bm{z}_{\star}+\bm{s}_{\star}$. By the definition of $\bm{s}$, we have $\mathcal{W}\bm{s}=\Gamma_{\gamma\alpha p}(\mathcal{W}\Pi_{\Omega}(\bm{z}_{\star}+\bm{s}_{\star}-\bm{z}))$, thus
@@
Combing all the pieces we have $\big{\|}\mathcal{W}\Pi_{\Omega}\bm{s}_{\star}-\mathcal{W}\bm{s}\big{\|}_{\infty}\leq 2\left\lVert\mathcal{W}\Pi_{\Omega}(\bm{z}_{\star}-\bm{z})\right\lVert_{\infty}$. ∎
Lemma 7.
@@
$\left\lVert(\widetilde{{\bm{R}}}\widetilde{{\bm{Q}}}^{-*}-\bm{R}_{\star})\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}$ $\leq\left\lVert(\widetilde{{\bm{R}}}\widetilde{{\bm{Q}}}^{-*}-\bm{R}_{\star})\bm{\Sigma}_{\star}^{1/2}\bm{\Sigma}_{\star}^{-1}\right\lVert_{\mathrm{F}}$ (14)
$\leq\left\lVert\bm{\Sigma}_{\star}^{-1}\right\lVert_{2}\mathrm{dist}\left(\widetilde{{\bm{L}}},\widetilde{{\bm{R}}};\bm{L}_{\star},\bm{R}_{\star}\right)\leq\varepsilon_{0}.$ (15)
@@
$\frac{C}{\big{\|}\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\big{\|}_{2}}\geq\frac{C}{(1+\varepsilon_{0})\big{\|}\widetilde{{\bm{L}}}_{i,:}\widetilde{{\bm{Q}}}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}}\geq\frac{\big{\|}[\bm{L}_{\star}]_{i,:}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}}{\big{\|}\widetilde{{\bm{L}}}_{i,:}\widetilde{{\bm{Q}}}\bm{\Sigma}_{\star}^{1/2}\big{\|}_{2}},$ (16)
@@
For any vectors $\bm{u},\bm{v}\in\mathbb{R}^{n}$, it holds that $\left\lVert\min(1,\lambda)\bm{u}-\bm{v}\right\lVert_{2}\leq\left\lVert\bm{u}-\bm{v}\right\lVert_{2}$ provided $\lambda\geq\frac{\left\lVert\bm{v}\right\lVert_{2}}{\left\lVert\bm{u}\right\lVert_{2}}$ (see [36, Claim 5]). Recalling that ${\bm{L}}_{i,:}=\min\left(1,\frac{C}{\big{\|}\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\big{\|}_{2}}\right)\widetilde{{\bm{L}}}_{i,:}$, thus by Equation 16 we have
@@
and similarly $\left\lVert\left({\bm{R}}_{i,:}\widetilde{{\bm{Q}}}^{-*}-[\bm{R}_{\star}]_{i,:}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2}\leq\left\lVert\left(\widetilde{{\bm{R}}}_{i,:}\widetilde{{\bm{Q}}}^{-*}-[\bm{R}_{\star}]_{i,:}\right)\bm{\Sigma}_{\star}^{1/2}\right\lVert_{2}^{2}$. Then, by definition of the metric we have the local non-expansiveness of $\mathcal{P}_{\mathcal{C}}$, i.e.,
@@
$\leq\mathrm{dist}^{2}\left(\widetilde{{\bm{L}}},\widetilde{{\bm{R}}};\bm{L}_{\star},\bm{R}_{\star}\right).$
Noticing that $\|{\bm{L}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{1/2}\|_{2}=\|\min(1,\frac{C}{\|\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{\frac{1}{2}}\|_{2}})\widetilde{{\bm{L}}}_{i,:}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{1/2}\|_{2}\leq C$, $\forall i\in[n_{1}]$, thus $\|{\bm{L}}{\bm{R}}^{*}\|_{2,\infty}\leq\|{\bm{L}}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{1/2}\|_{2,\infty}\|{\bm{R}}(\widetilde{{\bm{R}}}^{*}\widetilde{{\bm{R}}})^{-1/2}\|_{2}\leq C$. Similarly, we have $\|{\bm{R}}{\bm{L}}^{*}\|_{2,\infty}\leq C$. ∎
5.2 Proof of Proposition 1
Proof.
For $d_{k}\leq\varepsilon_{0}\sigma^{\star}_{r}$, following Equation 14, we have $\left\lVert\bm{\Delta}_{{\bm{L}}_{k}}\bm{\Sigma}_{\star}^{-1/2}\right\lVert_{2}\leq\varepsilon_{0},$ thus